OpenStudy (anonymous):

Solve by the substitution method 2x-5y=-19 7x+83=y I need help with the solution

OpenStudy (anonymous):

From the second equation you know y. Put that expression back into the first equation for y and solve for x. Then use that value for x in the second equation to solve for y.

jhonyy9 (jhonyy9):

2x-5(7x+83)=-19 2x-35x-415=-19 -33x=396 x= 396/-33 x= -132/11 7(-132/11)+83=y -924+913=11y -11=11y y= -1 x= -132/11 = -12 y= -1